Coolant = 20°c T h 500 W/m2 K Chip ke=50 W/m K q=107 W/m3


Coolant Chip k-50W/m.K q=10 W/m2 T 20°c oo h=500 W/m2 K HI4 3 Substrate, k5.Wm-K H= 12 mm L-27 mm
A simplified representation for cooling in very large- scale integration (VLSI) of microelectronics is shown in the sketch below. A silicon chip is mounted in a dielectric substrate, and one surface of the system is convectively cooled, while the reminding surfaces are well insulated from the surrounding. The problem is rendered two dimensional by assuming the system to be very large in the direction perpendicular to the paper. Under steady state (1) and transient (2) operation electronic power dissipation in the chip provides for uniform volumetric heating at a rate of q. However, the heating rate is limited by the restrictions on the maximum temperature that the chip is allowed to achieve. Cooling information is T-20 °C and h=500 W/m2.K the conductive heat transfer of chip is kc 50 W/m.K, and q = 107 WIms, the heat conduction coefficient for the substrate is ks 5 W/m.K For the condition shown on the sketch, will the maximum temperature inthe chips exceed 85 °C, the maximum allowable operating temperature set by industry standards? A grid spacing of 1.5 mm is suggested. Find the temperature distribution in the chips and substrate (1) Steady state two dimensions (2) Transient only for three successive nodes from the center of chip in two dimensions (a = 89.2 x 10-6 Plot the temperatures of this three successive nodes versus time within 5 min of chip's operation. Assumption: Constant Properties Thermal heat transfer coefficients are constant .Radiation heat trans fer negligible
Coolant = 20°c T h 500 W/m2 K Chip ke=50 W/m K q=107 W/m3
Coolant Chip k-50W/m.K q=10 W/m2 T 20°c oo h=500 W/m2 K HI4 3 Substrate, k5.Wm-K H= 12 mm L-27 mm
Answer #1

SOLUTION:

Perform an energy balance on the chip q(H/4) (H4) T h Tmax = 2kc 107 W/m3 (0.003 m)10 W/m2 (0.003 m) ,C = 8o 9C 2 x 50 W/mhAx 2 + hAx T (1) Ti+T T6 - 2.30T1 T2 + T6 = -6.00 Node 2 (2) TI 3.3T2 T3 + T7 = -6.00 Node 3 Ts - T3 kgAx Тз — Тз kgAy Та -matrix inversion method, finite-difference equations is written in matrix notation, [A][T] = [C] 10 0 0 0 0 0 0 0 0 0 0 0 0 -